CAT Chat Minutes

April 8, 2002

 

Information and Follow-up Items

Last Saturday night, our shutter closed for no apparent reason. During that time, the shutter permit was not taken away, and we lost beam all evening. What caused this error?

It was observed that many beamline shutters were closed on Sunday morning April 7 at 0335. The source of the problem was traced to a transient in the BTS:BX magnet power supply. This is the magnet in the booster-to-storage ring transport line which is responsible for directing the extracted booster beam either to the ring or to a beam dump. The access control system (ACIS) monitors the magnetic field in this magnet, and responded to the transient by very briefly removing beamline shutter permit, causing most shutters to close. Twenty nine shutters were affected, however only 22 had been re-opened one hour later, so seven shutters were closed for an extended period. The ultimate source of the problem was associated with the BTS:BX controls. Because of the intermittent action, no one was aware that shutter permit have been automatically removed, and no one, from the facility, was aware that the shutter was closed for any reason, other than user intervention. Efforts to improve communication to the staff and users following events of this nature will continue.

During the high current tests on April 1, unexpected changes in the injection pattern were observed. This happened every 30 seconds with large variations of intensity in our current monitor. What took place at that time?

Because of operational problems, the injector did not become available until after 10:30 AM. This was compounded by an operational error which caused a beam dump at 11:02. Once 102 mA was finally attained at 11:09, fully three hours behind schedule, the users were polled with regard to when 115 mA should be attempted, and the consensus was to try it as soon as possible. It would be accurate to characterize the high current study as very disappointing from an operations standpoint. All efforts will be made to assure a succesful high current study during the next run on Monday, June 24, from 0800 to 1600.

User Operations

Glenn Decker presented the seven-day storage ring current and beamline operations history. Glenn indicated availability is ~ 99.43% for the week. The mean time between failures has been 119.95 hours. One fault occurred caused by a beam motion probably due to the orbit correction.

Glen also presented the overall run of operating cycle 1-02. Overall availability is 97.22% with MTBF of 31.80 hours. Hours scheduled thus far are 1505 with 1463 being delivered.

A 48-hour machine intervention/studies period will start on Monday, April 15th beginning at 08:00 hours.

 

General Information

Tony displayed viewgraphs and discussed with the members of the group the reasons and some of the details of the reorganization of the APS structure. Apart from budget reasons, the reorganization was made to more effectively position the APS organization in providing increased user support. As an example, similar expertise groups were combined to increase effectiveness and efficiency within the organization. In the interim, all current responsibilities and contact personnel remain the same, and changes will be clearly defined and communicated to the users. The CAT Operations and support organization will continue to evolve, as plans are refined during the preparation of a response to the issues raised from the recent DOE review and as a result of the facility staff and user interactions during the upcoming retreat.